OpenStudy (yacoub1993):

HELP!!! Solve the equation. Check for extraneous solutions. |4x+3|=9+2x x= _________ or _________

OpenStudy (paki):

you will get 2 equations here... 4x-3=9+2x and 4x+3=9+2x solve for "x"...

OpenStudy (yacoub1993):

4x-2x=9+3 2x=12 x=12/2 x=6

OpenStudy (yacoub1993):

4x+3=9+2x 4x-2x=9-3 2x=6 x=3

OpenStudy (yacoub1993):

is that correct @paki

OpenStudy (paki):

yeah... correct..

OpenStudy (dumbcow):

absolute value equations must account for both positive and negative cases 4x+3 = 9+2x and 4x+3 = -(9+2x)